OpenStudy (anonymous):

Given: a and B are first-quadrant angles, sin a=5/13, and cosB= 3/5. evaluate sin(a-B). a.) -33/65 b.) 33/65 c.) 63/65

OpenStudy (ash2326):

Use this here \[\sin (A-B)=\sin A\times \cos B-\sin B\times \cos A\]

OpenStudy (ash2326):

sin a=5/13, and cosB= 3/5

OpenStudy (anonymous):

I tried this but the book is asking me to subtitute pi/2 for a

OpenStudy (anonymous):

and I know I have to solve for sin B and cos A

OpenStudy (ash2326):

\[\sin (B)=\cos (\frac {\pi}{2} -B)\] \[\cos A=\sin (\frac{\pi}{2}-{A})\] you can use \[\sin^2 x+\cos^2 x=1\] to figure out sin B and cos A

OpenStudy (anonymous):

ok ill try this

OpenStudy (anonymous):

cos a= 12/13

OpenStudy (ash2326):

correct :)